OpenStudy (anonymous):

X^2 + 10 what is the x intercept? I got sort of -10 which is an imaginary number I think

OpenStudy (anonymous):

x-intercept means y = 0. as y = x^2 +10 => x^2 +10 = 0 x^2 = -10 taking square root on both sides we get an imagnary number therefore the given equation doesnt intersect the x-axis

OpenStudy (unklerhaukus):

x^2 = -10 x = \(\sqrt{-10} \) x = \(\sqrt{-1\times10}\) x = \(\sqrt{10}\sqrt{-1}\) x = \(\pm\sqrt{10}i\)